# Multipoint Schwarz–Pick Lemma for quaternionic case

### 15 Dec 2023 v1

@@ 1.1. Slice regular functions
$(f\ast g)(q):=\sum_{m=0}^{\infty}q^{m}\left(\sum_{\ell=0}^{m}a_{m-\ell}b_{\ell}\right).$
This power series still converge in $B(0,R)$. Since the product in $\mathbb{H}$ is associative, then also the $\ast$-product is associative; however, it is clearly noncommutative.
@@
A function $f:\Omega\longrightarrow\mathbb{H}$ is slice preserving if for any $I\in\mathbb{S}$ it is $f(\Omega_{I})\subseteq\mathbb{C}_{I}$. It is one slice preserving if there exists $I\in\mathbb{S}$ so that $f$ is $\mathbb{C}_{I}$-preserving: $f(\Omega_{I})\subseteq\mathbb{C}_{I}$.
@@ 2. The hyperbolic difference quotient and the three-points Schwarz–Pick Lemma
Then $f_{p}^{\star}:\mathbb{B}\longrightarrow\mathbb{H}$ is a slice regular function as a $\ast$-product of regular functions: the last term is not singular since $f$ has values in the unit ball. In particular there exists $p_{0}\in\mathbb{B}$, depending on $f$ and $p$, such that $f^{\star}_{p}(p)$, called the hyperbolic derivative of $f$ at $p$, writes as
(7) $f^{h}(p):=f_{p}^{\star}(p)=(1-|p|^{2})\partial_{C}f(p)(1-\overline{f(p)}f(p_{0}))^{-1}.$
Note that $f_{p}^{\star}$ can also be written as
@@
(8) $\mathcal{M}_{f(p)}\bullet f=(f-f(p))\ast(1-\overline{f(p)}\ast f)^{-\ast}=(1-f\overline{f(p)})^{-\ast}\ast(f-f(p)).$
The quaternionic Schwarz–Pick Lemma
@@
The result is therefore not in general as in the complex case, where in the left hand side we just have that the pullback of the pseudo-hyperbolic distance with respect to an holomorphic function is less or equal than the pseudo-hyperbolic distance itself. Here the two maps $\mathcal{T}_{p}$ and $\widetilde{\mathcal{T}}_{f,p}$ appearing in the two sides are different!
@@
Three-points Schwarz–Pick Lemma
Now, let return to the hyperbolic difference quotient $f^{\star}_{p}$: we can apply the previous results obtaining a Schwarz–Pick type statement wich now involves three points of the unit ball.
@@
and the factor has clearly modulus one. ∎
@@
Let $c_{0}\in\mathbb{B}$ and $10<r_{0}<$ 1. Let consider the pseudo-hyperbolic ball $B_{\rho}(c_{0},r_{0}):=\{q\in\mathbb{B}:\rho_{\mathbb{B}}(q,c_{0})<r_{0}\}$. This coincides with the Euclidean ball $B(c_{1},r_{1}):=\{q\in\mathbb{H}:|q-c_{1}|<r_{1}\}$ where
@@
(17) $\left|f^{h}(q)\right|\leq\dfrac{\left|\partial_{C}f(0)\right|+\frac{2|q_{0}|}{1+|q_{0}|^{2}}}{1+\left|\partial_{C}f(0)\right|\frac{2|q_{0}|}{1+|q_{0}|^{2}}}.$
@@
$f^{h}(0)=f^{\star}_{0}(0)=(\mathcal{M}_{0}^{-\ast}\ast\mathcal{M}_{0}\bullet f)(0)=\partial_{C}f(0).$
@@
Distortion estimates for the class $\mathcal{B}_{\alpha}$
Let consider $\alpha\in[0,1)$ and let $\mathcal{B}_{\alpha}$ be the class of slice regular functions $f:\mathbb{B}\longrightarrow\mathbb{B}$ such that $f(0)=0$ and $\partial_{C}f(0)=\alpha$. Let us observe that in this case the Cullen derivative at zero coincides with the hyperbolic derivative $f^{h}(0)$ by the definition of the latter. For such functions we have the following estimates.
@@
Now, by Lemma 3.4 the inequality (19) says that for any $q\in\mathbb{B}$, $f^{h}(q)$ lies on an Euclidean ball contained in $\mathbb{B}$ which is symmetric with respect to $\mathbb{R}$ since $\alpha$ is real: the intersections with the real line of such ball are given by (14) where $c_{0}=\alpha$ and $r_{0}=2|q|/(1+|q|^{2})$:
@@ 4. Iterated hyperbolic difference quotients and the multipoint Schwarz–Pick Lemma
$\rho_{\mathbb{B}}\left(f^{\{n\}}_{p_{1},\ldots,p_{n}}(p),f^{\{n\}}_{p_{1},\ldots,p_{n}}\left(\widetilde{\mathcal{T}}_{f^{\{n\}}_{p_{1},\ldots,p_{n}},p}(q)\right)\right)\leq\rho_{\mathbb{B}}(p,\mathcal{T}_{p}(q)),$
@@
$\widetilde{\mathcal{T}}_{f^{\{n\}}_{p_{1},\ldots,p_{n}},p}(q):=h(q)^{-1}qh(q),\quad h(q)=1-(f^{\{n\}}_{p_{1},\ldots,p_{n}}(p)\ast(f^{\{n\}}_{p_{1},\ldots,p_{n}})^{c})(q).$
Let us observe that in the case $q\in\mathbb{R}\cap\mathbb{B}=(-1,1)$ the maps $\mathcal{T}_{p}$ and $\widetilde{\mathcal{T}}_{f^{\{n\}}_{p_{1},\ldots,p_{n}},p}$ act as identity on $q$ and we reduce to the classical Schwarz–Pick inequality in the pullback form:
@@ 5.1. Two points Nevanlinna–Pick interpolation
In the quaternionic frame we cannot expect an analogous result in general by the different expression of the Schwarz–Pick Lemma. However, if we assume that one of the point $p_{1},p_{2}$ is real, we can recover the same result.
@@
If there exists a function $f$ as required, then by the quaternionic version of Schwarz–Pick Lemma ([10, Thm 1.7]) it is
@@
Vice versa, let assume $\rho_{\mathbb{B}}(q,s)\leq\rho_{\mathbb{B}}(p,r)$. Then, by noticing that
@@
Let $r,p\in(-1,1)$, $r\neq p$ and $s,q\in{\mathbb{B}}$, such that the Pick matrix associated to the pairs $(r,p)$ and $(s,q)$ is positive semidefinite. If the Pick matrix is singular, then $Q:=M_{r}(p)^{-1}M_{s}(q)$ is unimodular and the interpolation problem
@@
By hypothesis $P\geq 0$, so there exists a solution in $\mathcal{SR}(\mathbb{B},\mathbb{B})$ for the interpolation problem $f(r)=s$ and $f(p)=q$. But since we assumed that $r$ and $p$ are real, then $f^{\star}_{r}(p)=(\mathcal{M}_{r}^{-\ast}\ast(\mathcal{M}_{f(r)}\bullet f))(p)$ can be computed as $M_{r}(p)^{-1}M_{f(r)}(f(p))=M_{r}(p)^{-1}M_{s}(q)=Q$.
@@
has the unique slice regular solution $g=\mathcal{M}_{-Q}\bullet(\mathcal{M}_{p}\ast h)$ for any $h\in\mathcal{SR}(\mathbb{B},\mathbb{B})\cup\partial\mathbb{B}$; fixed such a function $h$, now we can find $f$ so that
@@
So the condition for the existence of a solution for the interpolation problem is $(25-16\lambda^{2})\mu^{2}\leq 16-25\lambda^{2}$. In particular, if we fix $\lambda$ so that $1>|\lambda|>\frac{4}{5}$, we have no solutions for any choice of $\mu$; if otherwise $\lambda$ is fixed so that $0\leq|\lambda|\leq\frac{4}{5}$, then we have solutions once
@@
When equality holds, we have the extremal case of a unique (Möbius) solution, otherwise we have infinite solutions parameterized by $\mathcal{SR}(\mathbb{B},{\mathbb{B}})\cup\partial\mathbb{B}$.
For instance, let $\lambda=\frac{1}{4}$. Then we have solutions if and only if $|\mu|\leq\sqrt{\frac{77}{128}}$. When the inequality is strict, solutions are given (implicitely) by the parametric formula
@@ 5.2. Nevanlinna–Pick interpolation with three points
and  [^5] Here $\infty$ is just a conventional definition, it does not correspond to a pole of the ratio. For our purposes, we could have chosen any quaternion of norm $>1$. The other cases correspond to having as $Q_{1}^{2},Q_{1}^{3}$ two differents points on $\partial\mathbb{B}$, or a point inside $\mathbb{B}$ and the other one outside, or at least one of the two points not in $\overline{\mathbb{B}}$.
@@
Indeed $f^{\star}_{r_{1}}\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})\cup\partial\mathbb{B}$, hence $\overline{\mathbb{B}}\ni f^{\star}_{r_{1}}(r_{m})=\mathcal{M}_{r_{1}}^{-\ast}\ast(\mathcal{M}_{s_{1}}\bullet f)\mid_{r_{m}}$, $m=2,3$: since $r_{m}$ is real, the last term is equal to $\mathcal{M}_{r_{1}}(r_{m})^{-1}(\mathcal{M}_{s_{1}}\bullet f)(r_{m})=M_{r_{1}}(r_{m})^{-1}M_{s_{1}}(f(r_{m}))=M_{r_{1}}(r_{m})^{-1}M_{s_{1}}(s_{m})=Q_{1}^{m}$ for $m=2,3$. If $f$ is a Möbius trasformation, then $f^{\star}_{r_{1}}=f^{\{2\}}_{r_{1},r_{2}}\equiv\eta\in\partial\mathbb{B}$, so we have also $Q_{1}^{2}=Q_{1}^{3}=Q_{2}^{3}=\eta$. Otherwise, $f^{\star}_{r_{1}}$ has values in the open unit ball and $f^{\{2\}}_{r_{1},r_{2}}(r_{3})=(f^{\star}_{r_{1}})^{\star}_{r_{2}}(r_{3})$, which as above can be computed as $M_{r_{2}}^{-1}(r_{3})M_{f^{\star}_{r_{1}}(r_{2})}(f^{\star}_{r_{1}}(r_{3}))=M_{r_{2}}^{-1}(r_{3})M_{Q_{1}^{2}}(Q_{1}^{3})=Q_{2}^{3}$.
@@
Let assume $|Q_{2}^{3}|<1$. This implies by its definition that also $Q_{1}^{2},Q_{1}^{3}$ lie within the open unit ball. Hence by (24), if $f$ exists, then both $f^{\star}_{r_{1}}$ and $f^{\{2\}}_{r_{1},r_{2}}$ are slice regular maps from $\mathbb{B}$ to itself. Moreover we can define $f^{\{3\}}_{r_{1},r_{2},r_{3}}\in\mathcal{SR(\mathbb{B},\mathbb{B})}\cup\partial\mathbb{B}$. Now, let’s take $h$ arbitrarily in this latter set of functions and seek for a function $g_{1}$ whose hyperbolic difference quotient at $r_{3}$ is $h$ and attains the value $Q_{2}^{3}$ at $r_{3}$, i.e., the solution of
@@
We observe that the terms $Q_{\kappa}^{\ell}$ of the algorithm depend on the order we fix for the nodes of interpolation $r_{1},r_{2},r_{3}$, but the problem itself should not depend on that order. Hence, altough $Q_{2}^{3}$ could change, the property of being inside, outside or in the boundary of the unit ball $\mathbb{B}$ is preserved. Moreover, a permutation of nodes produces a different expression for the family of solutions (21) in the non-singular case and the unique solution (22), or (23), in the singular one: this is not a contradiction as those representations in terms of $\ast$-products and $\bullet$-actions are not unique.

### 25 Jan 2024 v2

# Multipoint Schwarz–Pick Lemma for the quaternionic case

@@ Abstract.
The two authors were partially supported by GNSAGA of INdAM and by PRIN Varietà complesse: geometria, topologia e analisi armonica
@@ 1.1. Slice regular functions
$(f\ast g)(q):=\sum_{m=0}^{+\infty}q^{m}\left(\sum_{\ell=0}^{m}a_{m-\ell}b_{\ell}\right).$
This power series still converges in $B(0,R)$. Since the product in $\mathbb{H}$ is associative, then also the $\ast$-product is associative; however, it is clearly noncommutative.
@@
A function $f:\Omega\longrightarrow\mathbb{H}$ is slice preserving if for any $I\in\mathbb{S}$ it is $f(\Omega_{I})\subseteq\mathbb{C}_{I}$. It is one slice preserving if there exists a unique $I\in\mathbb{S}$ so that $f$ is $\mathbb{C}_{I}$-preserving: $f(\Omega_{I})\subseteq\mathbb{C}_{I}$.
@@ 2. The hyperbolic difference quotient and the three-points Schwarz–Pick Lemma
Then $f_{p}^{\star}:\mathbb{B}\longrightarrow\mathbb{H}$ is a slice regular function as a $\ast$-product of regular functions: the last term is not singular since $f$ has values in the unit ball. In particular, its value in $q=p$ is called the hyperbolic derivative of $f$ in $p$ and it is denoted by $f^{h}(p)$. By evaluating the $\ast$-products, it is related to the regular derivative by the formula
(7) $f^{h}(p)=f_{p}^{\star}(p)=(1-|p|^{2})\partial_{C}f(p)(1-\overline{f(p)}f(p_{0}))^{-1},$
for some $p_{0}\in\mathbb{S}_{p}$ depending on $p$ and on the function $f$ itself. Note that $f_{p}^{\star}$ can also be written as
@@
(8) $\mathcal{M}_{f(p)}\bullet f=(f-f(p))\ast(1-\overline{f(p)}\ast f)^{-\ast}=(1-f\cdot\overline{f(p)})^{-\ast}\ast(f-f(p)).$
2.1. The quaternionic Schwarz–Pick Lemma
@@
The result is therefore not in general as in the complex case, where in the left hand side we just have that the pullback of the pseudo-hyperbolic distance with respect to an holomorphic function is less or equal than the pseudo-hyperbolic distance itself. Here the two maps $\mathcal{T}_{p}$ and $\widetilde{\mathcal{T}}_{f,p}$ appearing in the two sides are different.
@@
2.2. Three-points quaternionic Schwarz–Pick Lemma
Now, let us return to the hyperbolic difference quotient $f^{\star}_{p}$: we can apply the previous results obtaining a Schwarz–Pick type statement which now involves three points of the unit ball.
@@ 2.2. Three-points quaternionic Schwarz–Pick Lemma
and the factor $-(1-q\overline{p})^{-1}(1-p\overline{q})$ has clearly modulus one. ∎
@@
Let $c_{0}\in\mathbb{B}$ and $10<r_{0}<$ 1. Let us consider the pseudo-hyperbolic ball $B_{\rho}(c_{0},r_{0}):=\{q\in\mathbb{B}:\rho_{\mathbb{B}}(q,c_{0})<r_{0}\}$. This coincides with the Euclidean ball $B(c_{1},r_{1}):=\{q\in\mathbb{H}:|q-c_{1}|<r_{1}\}$ where
@@
(17) $\left|f^{h}(q_{0})\right|\leq\dfrac{\left|\partial_{C}f(0)\right|+\frac{2|q_{0}|}{1+|q_{0}|^{2}}}{1+\left|\partial_{C}f(0)\right|\frac{2|q_{0}|}{1+|q_{0}|^{2}}}.$
@@
$f^{h}(0)=f^{\star}_{0}(0)=(\mathcal{M}_{0}^{-\ast}\ast(\mathcal{M}_{0}\bullet f))(0)=q^{-1}f(q)\mid_{q=0}=\partial_{C}f(0).$
@@
3.1. Distortion estimates for the class $\boldsymbol{\mathcal{B}_{\alpha}}$
Let $\alpha\in[0,1)$ and let $\mathcal{B}_{\alpha}$ be the class of slice regular functions $f:\mathbb{B}\longrightarrow\mathbb{B}$ such that $f(0)=0$ and $\partial_{C}f(0)=\alpha$. Let us observe that in this case the Cullen derivative at zero coincides with the hyperbolic derivative $f^{h}(0)$ by the definition of the latter. For such functions we have the following estimates.
@@ 3.1. Distortion estimates for the class $\boldsymbol{\mathcal{B}_{\alpha}}$
Now, by Lemma 3.4 the inequality (19) says that for any $q\in\mathbb{B}$, $f^{h}(q)$ lies on a Euclidean ball contained in $\mathbb{B}$ which is symmetric with respect to $\mathbb{R}$ since $\alpha$ is real: the intersections with the real line of such ball are given by (14) where $c_{0}=\alpha$ and $r_{0}=2|q|/(1+|q|^{2})$:
@@ 4. Iterated hyperbolic difference quotients and the multipoint Schwarz–Pick Lemma
$\rho_{\mathbb{B}}\left(f^{\{n\}}_{p_{1},\ldots,p_{n}}(p),f^{\{n\}}_{p_{1},\ldots,p_{n}}\left(\widetilde{\mathcal{T}}_{(f^{\{n\}}_{p_{1},\ldots,p_{n}}),p}(q)\right)\right)\leq\rho_{\mathbb{B}}(p,\mathcal{T}_{p}(q)),$
@@
$\widetilde{\mathcal{T}}_{(f^{\{n\}}_{p_{1},\ldots,p_{n}}),p}(q):=h(q)^{-1}qh(q),\quad h(q)=1-(f^{\{n\}}_{p_{1},\ldots,p_{n}}(p)\ast(f^{\{n\}}_{p_{1},\ldots,p_{n}})^{c})(q).$
Let us observe that in the case $q\in\mathbb{R}\cap\mathbb{B}=(-1,1)$ the maps $\mathcal{T}_{p}$ and $\widetilde{\mathcal{T}}_{(f^{\{n\}}_{p_{1},\ldots,p_{n}}),p}$ act as identity on $q$ and we reduce to the classical Schwarz–Pick inequality in the pullback form:
@@ 5.1. Two points Nevanlinna–Pick interpolation
In the quaternionic framework we cannot expect an analogous result in general by the different expression of the Schwarz–Pick Lemma. However, if we assume that one of the point $p_{1},p_{2}$ is real, we can recover the same result.
@@
If there exists a function $f$ as required, then by (11) we recall that
@@
Vice versa, let us assume $\rho_{\mathbb{B}}(q,s)\leq\rho_{\mathbb{B}}(p,r)$. Then, by noticing that
@@
Let $r,p\in(-1,1)$, $r\neq p$ and $s,q\in{\mathbb{B}}$, such that the Pick matrix associated to the pairs $(r,p)$ and $(s,q)$ is positive semidefinite, or equivalently (20) holds. If the Pick matrix is singular, then $Q:=M_{r}(p)^{-1}M_{s}(q)$ is unimodular and the interpolation problem
@@
By hypothesis $P\geq 0$, hence by Theorem 5.1 there exists a solution in $\mathcal{SR}(\mathbb{B},\mathbb{B})$ for the interpolation problem $f(r)=s$ and $f(p)=q$. But since we assumed that $r$ and $p$ are real, then $f^{\star}_{r}(p)=(\mathcal{M}_{r}^{-\ast}\ast(\mathcal{M}_{f(r)}\bullet f))(p)$ can be computed as $M_{r}(p)^{-1}M_{f(r)}(f(p))=M_{r}(p)^{-1}M_{s}(q)=Q$.
@@
has the unique slice regular solution $g=\mathcal{M}_{-Q}\bullet(\mathcal{M}_{p}\ast h)$ for any $h\in\mathcal{SR}(\mathbb{B},\mathbb{B})\cup\partial\mathbb{B}$; once we fix such a function $h$, we can find $f$ so that
@@
So the condition for the existence of a solution for the interpolation problem $|Q|\leq 1$ in this case becomes $(25-16\lambda^{2})\mu^{2}\leq 16-25\lambda^{2}$. In particular, if we fix $\lambda$ so that $1>|\lambda|>\frac{4}{5}$, we have no solutions for any choice of $\mu$; if otherwise $\lambda$ is fixed so that $0\leq|\lambda|\leq\frac{4}{5}$, then we have solutions once
@@
When equality holds ($|Q|=1$), we have the extremal case of a unique (Möbius) solution, otherwise we have infinite solutions parameterized by $\mathcal{SR}(\mathbb{B},{\mathbb{B}})\cup\partial\mathbb{B}$.
For instance, let $\lambda=\frac{1}{4}$. Then we have solutions if and only if $|\mu|\leq\sqrt{\frac{77}{128}}$. When the inequality is strict, solutions are given (implicitly) by the parametric formula
@@ 5.2. Nevanlinna–Pick interpolation with three points
and  [^5] Here $\infty$ is just a conventional definition, it does not correspond to a pole of the ratio. For our purposes, we could have chosen any quaternion of norm $>1$. The other cases correspond to having as $Q_{1}^{2},Q_{1}^{3}$ two different points on $\partial\mathbb{B}$, or a point inside $\mathbb{B}$ and the other one outside, or at least one of the two points not in $\overline{\mathbb{B}}$.
@@
Indeed $f^{\star}_{r_{1}}\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})\cup\partial\mathbb{B}$, hence $\overline{\mathbb{B}}\ni f^{\star}_{r_{1}}(r_{m})=\mathcal{M}_{r_{1}}^{-\ast}\ast(\mathcal{M}_{s_{1}}\bullet f)\mid_{r_{m}}$, $m=2,3$: since $r_{m}$ is real, the last term is equal to $\mathcal{M}_{r_{1}}(r_{m})^{-1}(\mathcal{M}_{s_{1}}\bullet f)(r_{m})=M_{r_{1}}(r_{m})^{-1}M_{s_{1}}(f(r_{m}))=M_{r_{1}}(r_{m})^{-1}M_{s_{1}}(s_{m})=Q_{1}^{m}$ for $m=2,3$. If $f$ is a Möbius transformation, then $f^{\star}_{r_{1}}=f^{\{2\}}_{r_{1},r_{2}}\equiv\eta\in\partial\mathbb{B}$, so we have also $Q_{1}^{2}=Q_{1}^{3}=Q_{2}^{3}=\eta$. Otherwise, $f^{\star}_{r_{1}}$ has values in the open unit ball and $f^{\{2\}}_{r_{1},r_{2}}(r_{3})=(f^{\star}_{r_{1}})^{\star}_{r_{2}}(r_{3})$, which as above can be computed as $M_{r_{2}}^{-1}(r_{3})M_{f^{\star}_{r_{1}}(r_{2})}(f^{\star}_{r_{1}}(r_{3}))=M_{r_{2}}^{-1}(r_{3})M_{Q_{1}^{2}}(Q_{1}^{3})=Q_{2}^{3}$.
@@
Let us assume $|Q_{2}^{3}|<1$. This implies by its definition that also $Q_{1}^{2},Q_{1}^{3}$ lie within the open unit ball. Hence by (25), if $f$ exists, then both $f^{\star}_{r_{1}}$ and $f^{\{2\}}_{r_{1},r_{2}}$ are slice regular maps from $\mathbb{B}$ to itself. Moreover we can define $f^{\{3\}}_{r_{1},r_{2},r_{3}}\in\mathcal{SR(\mathbb{B},\mathbb{B})}\cup\partial\mathbb{B}$. Now, let’s take $h$ arbitrarily in this latter set of functions and seek for a function $g_{1}$ whose hyperbolic difference quotient at $r_{3}$ is $h$ and attains the value $Q_{2}^{3}$ at $r_{3}$, i.e., the solution of
@@
We observe that the terms $Q_{\kappa}^{\ell}$ of the algorithm depend on the order we fix for the nodes of interpolation $r_{1},r_{2},r_{3}$, but the problem itself should not depend on that order. Hence, although $Q_{2}^{3}$ could change, the property of being inside, outside or in the boundary of the unit ball $\mathbb{B}$ is preserved. Moreover, a permutation of nodes produces a different expression for the family of solutions (22) in the non-singular case and the unique solution (23), or (24), in the singular one: this is not a contradiction as those representations in terms of $\ast$-products and $\bullet$-actions are not unique.